Mr. Scott’s take Celebrex 300 mg IV STAT for grand mal seizures. Cerebrexx 75 mg/ml in a 10 Vialis available for administration.how many ml will the patient receive?


Mr. circle who underone a partial craniotomy dalantin100 mg IV every 8 hours how many millimeters will you prepare if the bottle available has 250mg in 5ML?


Timentin 3.1 g in 100 ML of D5W over 1hour is ordered fir s patient with sepsis. the drop factor is 60 gtt/ min. how many drop per minutes should the nurse give?



The physician orders a standard Z pack to be given over 5days the first day loading dose of 500mg, followed by 250 mg by for 4 days additional days. how many tablets are dispensed in the pack?



A client suffer from insomia. tha doctors ambien 5mg at bed bedtimeand the nurse can repeat the dose in one hours if needed. the nurse pharmacy send 10mg yablets how many tablets should the patient initially receive at bedtime?
Answer #1

●Given

Order:Celebrex

Dose:300mg

Route:IV

Frequency:Stat

Available:75mg/ml

When 1ml=75mg

Them 300mg=??ml

Cross multiply and divide

=300×1÷75

=300÷75

=4

Ans:4ml has to be administered

●Given

Order:Dalantin

Dise:100mg

Route:IV

Frequency :Three times a day

Available:250mg/5ml

When 250mg=5ml

Then 100mg=??ml

Cross multiply and divide

=100×5÷250

=500÷250

=2ml

Ans:2ml has to be administered

●Given

Order:Timentin

Dose:3.1g

Time:1hour (60minutes)

Drop factor:60gtt/mivolume:100ml

Flow rate=volume(ml)÷ Minutes ×drop factor

=100÷60×60

=1.67×60

=100

Ans:100drops/minute

●Given

The first day loading dose is 500mg=1 tablet

Then followed by 250mg/day for 4 days=4 tablet

So 5 tablets will be dispensed in the pack